OpenStudy (anonymous):

What is the integral of √x-1? I'm really behind my class and if someone could help explain this to me, I'd really appreciate it!

OpenStudy (amistre64):

18 more loki....come on!! you can do it ;)

OpenStudy (anonymous):

When you have a linear expression like,\[(ax+b)^{n}\]the solution will take the form,\[\int\limits_{}{}(ax+b)^{n}=\frac{1}{a}\frac{(ax+b)^{n+1}}{n+1}\]

OpenStudy (anonymous):

No need to waste time with a substitution like, u=x+1.

OpenStudy (anonymous):

hehe, amstre :)

OpenStudy (anonymous):

You can always check the result of your integration by differentiating the answer and comparing it with the integrand (the 'stuff' under the integral sign).

OpenStudy (anonymous):

Your case, the answer is\[\frac{2}{3}(x-1)^{3/2}+c\]

OpenStudy (anonymous):

I left out the constant on the general solution.

OpenStudy (anonymous):

I'm assuming your integral was \[\int\limits_{}{}\sqrt{x-1}dx\]

OpenStudy (anonymous):

Yes, thank you so much. That helps a ton. I apologize for being offline when I should have been on, as well. Things got out of control here on my end and I had to go help. Thank you!

OpenStudy (anonymous):

You're welcome :)